OpenStudy (anonymous):

check my answer please. Suppose that the functions p and q are defined as follows. p(x)=x^2+3 q(x)=sqrt(x+2) find (q*p)(2)=[?] and (p*q)(2)=[?] I got (q*p)(2)=sqrt 10 and (p*q)(2)=8

OpenStudy (anonymous):

\[p(x)=x^2+3\]\[q(x)=\sqrt{x+2}\] \[(qp)(x)=q(x) \times p(x)\] \[pq(x)=p(x) \times q(x)\] Since multiplication is commutative, I don't understand how you are getting different answers

OpenStudy (usukidoll):

this could be a composite function question like (q o p) (x) or (p o q)(x) ?

OpenStudy (anonymous):

I was thinking that too, let me check your answers @Kimes

OpenStudy (anonymous):

sorry yes (q o p) (x) or (p o q)(x) i the right format

OpenStudy (anonymous):

I think your answers are incorrect @Kimes , both of them

OpenStudy (anonymous):

well I followed the steps shown in my book, can you explain then

OpenStudy (anonymous):

Can you show me your steps, how you've attempted the question?

OpenStudy (anonymous):

(q o p) q(p(2)) q(2^2+3) q(7) sqrt 7+3= sqrt 10 (p o q) p(q(2)) p(sqrt 2+3) p(sqrt 5) (sqrt 5)^2+3=8

OpenStudy (anonymous):

I see the mistake now, let me ask you first what is your q(x)? \[q(x)=\sqrt{x+2}\] \[q(x)=\sqrt{x+3}\] Because in the question you've said sqrt{x+2} but in the solution you've used sqrt{x+3}

OpenStudy (anonymous):

ahh I made a silly mistake! thanks for catching it!

OpenStudy (anonymous):

You're welcome, your method of doing the question is perfectly fine, which is good as that means you've grasped the concept, only thing is you made a calculation mistake, that's not a big problem you just have to be careful